A 500g metal ball is brought to the moon. What is its weight on Earth? on the moon? What is its mass on Earth? on the moon?


1
Expert's answer
2021-12-17T11:49:47-0500

1. The weight is given as follows:


"W = mg"

where "m = 0.5kg" is the mass of the ball, "g" is the gravitational acceleration. On Earth: "g\approx 9.8m/s^2", on Moon: "g\approx 1.6m/s^2".

Thus, obtain weight on Earth:


"W_E = 0.5kg\cdot 9.8m/s^2 = 4.9N"

On Moon:


"W_M = 0.5kg\cdot 1.6m/s^2 = 0.8N"

2. The mass is always the same regardless of the planet.


Answer. Weight on Earth: 4.9N, weight on Moon: 0.8N, mass on Earth: 0.5kg, mass on Moon: 0.5kg.